OpenStudy (anonymous):

Determine the constants a and b given that the parabola y=ax^2+bx+1 passes through (-1,11) and (3,1).

OpenStudy (anonymous):

if you replace x by -1 you get \[11=a-b+1\] and so \[a-b=10\] likewise replace x by 3 and get \[1=9a+3b+1\] so \[9a+3b=0\] solve the system